PartnerObserving Log for 2007 Oct 06-07

Observers: J. Hill, V. Testa
Instrument Support: R. Bertram
Telescope Operator: A. Ceranski

005954 - 010215 - 5 bias frames

01:08 UT Open and watch sunset with UofA SEDS group. T=3degC, D=1.5 degC, Wind SSW at 7 m/sec, sky full of wispy clouds.

Sky flats with thin clouds everywhere (Vincenzo)

01:23 - SktFlatTestUspec (main mirror fail to focus execution continues)

012444 counts ~ 8500

01:25 - 5 flats Uspec rot=90 exp.time x 9.00
012634-013018 ~40000 - ~17000

01:31 - 5 flats Uspec rot=270 exp.time x 20.00
013319-013642 ~17000 - ~10000

01:37 - 5flats V rot=270 exp.time x 7.00
013901-014208 ~10000 - ~10000

Rotator Center and Pointing Tests (John)

Starting with the default collimation lookup table as sky is too bright for random pupils. Coma plus alot of astigmatism caused by actuator 140 being disabled.

014717 - 1 sec V on pointing star (not in pointing log)

start pointing log in file /home/telescope/100607PointingLogJH

015207 - 1 sec V on pointing star (paragraph boundaries here mean a new preset)
015250 - 10 sec V at PA=0
015502 - 10 sec V at PA=180

020013 - 10 sec V at PA=180
020157 - 10 sec V at PA=0
020311 - 1 sec V at PA=0

020515 - Vfastextra, sent Z4=2033
020704 - Vfastextra, no good pupils found (too much coma with astigmatism), sent Z7=+1000 manually
021010 - Vfastextra, couldn't believe this fits to the quite elliptical pupils. sent Z11=+500 manually

02:14 - Close dome for high humidity.

02:52 - Reopen and continue pointing tests. T=2.6degC, D=1degC. Wind gusting to 17 m/sec from SSW. Sky clearing.

030125 - Vfastextra, sent Z4 -3412 plus another Z11 +500 manually 030521 - Vfastextra, sent -1026 3897 1302 472 -75 -72
030730 - Vfastextra, sent -388 588 133 460 58 398
030948 - Vfastextra, sent -704 304 686 168 184 61
031141 - Vfastextra, sent -731 -254 -684 305 -20 -71
Note the trefoil residuals that lbcfpia doesn't know how to measure.

Now some round images for the rotator center measurement (near meridian)
031417 - 10 sec V at PA=0
031600 - 10 sec V at PA=180
031742 - 10 sec V at PA=0
031940 - 10 sec V at PA=180

032150 - Vfastextra, sent -708 694 -75 -7 -117 -115
032354 - Vfastextra, sent -476 -405 548 388 263 -72 (note rotation jump between these two images)

Now collecting a few pointing stars to make a more careful check around the sky.

033302 - pts_002 0017-01
033412 - V5sec at PA=0 using the same pointing = no preset
033542 - V5sec at PA=180
033721 - V5sec at PA=0 (pointing differs from 033302 by 1.4 arcsec)

034048 - pts_004 0057+21

034326 - pts_006 0119+45 (pointing differs by 1.5 arcsec from the next three presets to the same star)

034439 - pts_006 preset again on the same star

034529 - pts_006 preset again on the same star

034618 - pts_006 preset again on the same star

035143 - pts_069 2118-17 [1130,3011]

035453 - pts_068 2055+07 [981,2997]

035816 - pts_067 2030+27 [970, 2857]

040252 - pts_065 19:35+46 [1007,2957]

Preliminary conclusion is that this pointing model with zeropoint adjustment earlier in the week is usable for SDT, so we are handing over to Vincenzo to do some science in a very clear sky. The major uncertainty in the pointing just now is the inconsistency of our determination of the rotator center. I had trouble running rotcenter.cl (cannot open uparm$pipe9660 for writing?), so I've been analyzing the PA pairs above by hand with DS9.

The first conclusion is that the "old" rotator center (pre 22-Sep-2007) is clearly better than the new one. Therefore as of 08:00 UT, I have restored the LBC header files to the CRPIX values that correspond to a rotator center of (1043,2951). These CRPIX values have not been applied to the images taken between 22-Sep and 08:00 UT on 07-Oct.

The second conclusion is that (based on 4 stars), the present pointing error is about 20 arcsec rms. This is indeed good enough for the present SDT observations. In case we get a night of horrible seeing or heavy cirrus, we'll commandeer some more SDT time to collect better pointing data. Otherwise, more pointing data isn't much use until we solve the jitter from image to image, and understand our rotator center measurements better.

SDT Science Observations (Vincenzo)

1st target: Nonino CL0016_C_1
042034 - Vfastextra
041213 - Vfastextra

04:13 - CL0016_C_1 240s x 5 (Airmass: 1.18, filter: U)

041442: seeing = 1.7"
041935: seeing = "
042436: seeing = 1.5"
042924: seeing = "
043418:

2nd target: Ryan Oct 15 Deg Epoch 1:
Image Analysis:044246 sent -2027 0 0 0 0 0
Image Analysis: 044427 sent 53 318 650 -63 -29 88

04:46 - Oct 15 deg epoch 1 240s x 2 (Airmass: 1.33, filter: V)

044708: seeing = 1.4"
045231: seeing = 1.4"

04:59 3rd target: Oct 10 deg epoch 1
Image Analysis: 045916 sent -162 18 -565 0 82 -19

05:00 - Oct 10 deg epoch 1 240x2 (Airmass: 1.35, filter:V)

050125: seeing = 1.35"
050651: seeing = 1.45"

05:15 - SA113_239 8pos Vgr (Airmass: 1.25, filter; V,g,r)

051635-053530 Humidity rises, 92.6% @05:30

05:36 - Move to Ryan target Oct 5 deg epoch 1
Image Analysis: 054059 - No GOOD pupils found... check input image!
Preset 10 arcmin away and retry - Image Analysis: 054539 sent 744 -324 506 386 -84 -113

05:48 - Close again for rising humidity. Right shutter door blew closed. Left shutter door overcurrented the drives fighting the wind. Left door successfully closed 5 minutes later after Aaron reset the drives.

08:22 - Stopped the Pointing Log file

10:34 - Reopen and continue science observations. T=0.3degC. D=-4.5degC, Humidity 83% and rapidly falling. Wind 10 m/sec from SSW.

10:39 - 1st target, 2nd epoch: Ryan Oct 15 deg epoch 2
Image analysis: 104022 sent -444 -77 -557 -601 -116 -152

104300: seeing = 1.4"
104821: seeing = 1.4"

10:53 - Oct 10 deg epoch 2
Image Analysis: 105457 No GOOD pupils found... check input image!

10:55 - Telescope stops tracking apparently because HBS oil shut off. (JMH changed the oil temperature a little before this.) Lack of oil shuts off building drive -- we suppose. Wind blows building into telescope and trips MCR relay. Why didn't building brakes set? An alternate theory is that wind blew the building into the telescope, but that doesn't explain why oil was off. Lost 20 minutes to HBS.

11:15 - While recovering from above, AZREADY won't turn on Az drive because wind is blowing telescope. Can turn on Az servo with individual commands, but then AZREADY turns it off again. Telescope Azimuth brakes won't set. Close shutter doors to reduce sail area. Turn on Az drive and reopen shutter doors. Lost 20 minutes to MCSPU.

11:30 - LBCFPIA won't recognize the funny shaped pupils (which we think are caused by Act 140) at low elevation. Changing fields didn't help. Increasing exposure time didn't help. Clearing active optics didn't help. Finally had to go back to higher elevation and try again. The problem was not that we got bad fits to the pupils but that it could not identify the pupils at all. Lost 30 minutes to LBCFPIA.

Image Analysis: 112800 No GOOD pupils found... check input image!
Image Analysis: 112932 (32 sec exposure) No GOOD pupils found... check input image!

Image Analysis: 113201 (new field) -669 -1774 3106 -281 -893 -211
Image Analysis: 11334 No GOOD pupils found... check input image!
11:35 - Cleared "Active Optics" on PSF panel
11:39 - Image Analysis: 113926 -6706 0 0 0 0 0
11:41 - Image Analysis: 114133 -394 688 1188 535 -91 -63
11:43 - Image Analysis: 114306 No GOOD pupils found... check input image!

11:44 - Ryan Oct 10 deg epoch 2
114538: (quite horrible astigmatism on guiding images) aborted after the first, no images saved at all

[Note added later: some of this may have been the mirror oscillation that we noticed the following night. JMH]

cleared wavefront coefficients (C00)

11:50 - Image Analysis: 115115 -622 2879 1155 2662 -202 621
11:52 - Image Analysis: 115256 No GOOD pupils found... check input image! Compensate Z7 = -2000
11:55 - Image Analysis: 115531 No GOOD pupils found... check input image!

11:58 - Moved to a standard field at higher elevation: SA95 @ 55deg, to check if Vfastextra works at smaller airmass
12:00 - Image Analysis: 120023 -891 1705 672 228 391 512
12:02 - Image Analysis: 120219 -569 854 -328 134 195 177

12:06 - Move to UMajorII_V_1 (Sand) (Airmass= 1.33)
12:09 - Image Analysis: 121005 169 -265 806 57 -120 15
121222: seeing = 1.44 background: 13000
121810: seeing = 1.40 background: 27000
12:25 - aborted OB. Wait for skyflats

Twilight is upon us ..... we've almost forgotten what morning twilight is.

Morning Sky Flats (Vincenzo)

12:28 - go to flat field pointing and start with SkyFlatTestV (bias ~ 670)
123422: counts/sec = ~800 (1440-670)

12:36 - SkyFlatVrot90_5 (exp.time = 20.00s) expect ~ 16000-20000

123718: counts = 26300
123819: " = 27500
123921: " = 34000
124023: " = 41000
124136: " = 63000 (!)

12:43 - SkyFlatVrot270_5 (ex. time = 10.00) expect ~ 35000-45000

124426: counts = 52000
124517: " = saturated
124606: " = saturated
reset exp time = 4.0 s
124713: " = 47000
124807: " = 59032 (saturated)

12:49 - redo OB at exp. time = 1.0 s

124950: counts = 23000
125030: " = 27000
125112: " = 32000
125152: " = 39000
125239: " = 48000

12:53 - SkyFlatTestUspec: counts/sec = 17000

12:56 - SkyFlatUspec (exp. time = 1.0 s) expect ~ 20000-40000

125650: counts = 22500
125730: " = 33000
125816: " = 39000
125902: " = 48000
125942: " = 55500

12:49 - redo OB at exp. time = 1.0 s

124950: counts = 23000
125030: " = 27000
125112: " = 32000
125152: " = 39000
125239: " = 48000

12:53 - SkyFlatTestUspec: counts/sec = 17000

12:56 - SkyFlatUspec (exp. time = 1.0 s) expect ~ 20000-40000

125650: counts = 22500
125730: " = 33000
125816: " = 39000
125902: " = 48000
125942: " = 55500

13:01 - Close dome because it is getting light outside. T=-0.2degC, D=-3.7degC, Wind 9 m/sec from South.

25 biases.

-- JohnHill - 06 Oct 2007
Topic revision: r7 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback